No. Of hyperconjugated H-atom in

(C2H5)2-CH=CH-(C2H5)2
1
Expert's answer
2019-08-22T03:22:19-0400

(C2H5)2-CH=CH-(C2H5)2 does not exist, it is a mistake.

If it had to be (C2H5)-CH=CH-(C2H5), then there are 4 hyperconjugable hydrogen atoms.

If it had to be (C2H5)2-C=C-(C2H5)2, then there are 8 hyperconjugable hydrogen atoms.